OpenStudy (anonymous):

Mikayla has 15 coins made up of all dimes and quarters worth a total of $2.85. d = number of dimes, q = number of quarters Which system of equations represents this word problem? A. d + q = 15 0.10d + 0.25q = 2.85 B. d + q = 2.85 0.10d + 0.25q = 15 C. d + q = 15 10d + 25q = 2.85 D. d + q = 2.85 10d + 25q = 15

OpenStudy (anonymous):

D+q=15 ( the sum of both must be 15) .10d+.25q=2.85 ( the value of which coin time their quantity gives you the total value) Latter A ;)